kimperez123:

Bob fenced in his backyard. The perimeter of the yard is 22 feet, and the length of his yard is 5 feet. Use the perimeter formula to find the width of the rectangular yard in inches: P = 2L + 2W.

jhonyy9:

given that the perimeter P=22 feet and the length L = 5 feet and the formula P = 2L +2W just substitute the dates what you know in their place and calcule the width

kimperez123:

i still dont understand

jhonyy9:

given formula P = 2L +2W and you know P=2 feet and L=5 feet so just substitute these inside formula so what you get in this way

jhonyy9:

sorry P=22 feet

jhonyy9:

given that P=22 so substitute in place of P the 22 and you know that the length l=5 feet so substitite in place of L the 5 and what will get ?

jhonyy9:

P= 2L +2W 22 = 2*5 +2W yes ?

jhonyy9:

ok. ?

kimperez123:

so for the 2*5 is multiply or divide?

jhonyy9:

there is 2 time 5

kimperez123:

so it would be 6in?

jhonyy9:

how you get it ?

kimperez123:

by mutiplying

jhonyy9:

step by step please

kimperez123:

The formula f equals c over lambda, where f = frequency, c = wave speed, and λ = wavelength, is used to calculate frequency. Solve this formula for c.

jhonyy9:

22 = 2*5 +2W yes ?

kimperez123:

22=10=2w -2w=10-22 -2w=-12 w=6

jhonyy9:

yes right

kimperez123:

can you help with another question?

jhonyy9:

i will try post it on the openquations column please

kimperez123:

ok

kimperez123:

wait how do i do this?

kimperez123:

new at this

kimperez123:

Solve the following equation for x: 2x + 5y = 8.

jhonyy9:

you dont know the y value ?

jhonyy9:

like a first step substract from both sides 5y

jhonyy9:

in this way what will get ?

jhonyy9:

please collaborate

kimperez123:

2x=8-5y x=4-5/2y

jhonyy9:

yes but correctly is x = 4-(5/2)y

kimperez123:

thank you

kimperez123:

Solve the following equation for x: 2x − 3y = 6.

kimperez123:

?????